Solve each system by the substitution method.
3x+y=2
-x-3y=6

Answer by checkley71(8403)

3X+Y=2 OR Y=-3X+2 NOW SUBSTITUTE (-3X+2) FOR Y IN THE OTHER EQUATION:
-X-3(-3X+2)=6
-X+9X-6=6
8X=6+6
8X=12
X=12/8
X=1.5 ANSWER.
Y=-3*1.5+2
Y=-4.5+2
Y=-2.5 ANSWER.
PROOF
-1.5-3*-2.5=6
-1.5+7.5=6
6=6